OpenStudy (anonymous):

a line with a slope of 2 passes through (3,9). what is the equation of this line?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ 3}}\quad ,&{\color{blue}{ 9}})\quad % (c,d) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies 2 \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

point-slope form

OpenStudy (candy13106):

Y-9=2(x-3) do you understand how I got that